Mechanics

A particle moving With a velocny of (4i - j) m/s strikes a fixed smooth wall and finally moves with a velocity of (3i + 2j) m/s. The coefficient of restitution between the wall and the particte in the collision is:

Sarthak Khandelwal
8 Years agoGrade

1 Answer

venkat
ApprovedApproved Tutor Answer8 Years ago
 Given,
initial velocity of the ball u1=​\sqrt{}17m/s
initial velocity of wall is u2=0m/s
final velocity of the ball v1​=\sqrt{}13m/s
final velocity of the ball v2​=0m/s
Coefficient of restitution 
          e=(v2-v1)/(u1-u​1)
          e=\sqrt{}(13/17)